# Diffusion Models, Image Super-Resolution And Everything: A Survey

### v2

@@ Abstract
Image Super-Resolution, Diffusion Models, Generative Models, Survey.
@@ I Introduction
In the ever-evolving field of computer vision, the task of image Super-Resolution (SR) – enhancing Low-Resolution (LR) images into High-Resolution (HR) counterparts – stands as a longstanding yet still perplexing challenge, owing to its inherently ill-posed nature: Multiple HR reconstructions are plausible for any given LR image, differing in aspects such as brightness and color [1]. The applications span a broad spectrum, from enhancing everyday photography to refining satellite [2] and medical images [3], all driven forward by recent rapid advancements in deep learning.
The advent of Diffusion Models (DMs) marks a significant shift in generative image generation tasks, including image SR, challenging the long-standing dominance of Generative Adversarial Networks (GANs) [4, 5, 6, 7]. Despite the notable achievements of previous generative models, each comes with its set of limitations. For example, the computational demands of Autoregressive models often outweigh their utility in generating HR images [8, 9, 10], while Normalizing Flows (NFs) and Variational Autoencoders (VAEs) struggle to match the quality expectations. GANs, although powerful, necessitate careful regularization and optimization strategies to mitigate issues like optimization instability and mode collapse [11].
@@
In this context, DMs have not only contributed significantly to SR but have also heralded a new epoch in generative AI. Their pivotal roles in applications such as Dall-E and Stable Diffusion demonstrate that DMs have surpassed GANs in various aspects, redefining the generative models landscape [12, 13, 14]. Their capability to generate high-fidelity, detailed images from LR inputs has shown immense promise in SR, closely aligning with the qualitative judgments of human evaluators and offering a novel approach to image quality enhancement [15]. In other words, human raters perceive SR images generated by DMs as more realistic than those generated by GANs.
@@
However, as the volume of research on DMs continues to expand, it is becoming more challenging, particularly for those new to the field, to stay updated on the latest developments. DMs diverge fundamentally from their generative predecessors, posing new challenges while addressing the limitations of earlier models. Amidst this rapid expansion, identifying coherent trends and potential research directions can be challenging. This survey aims to demystify the current landscape of DMs in SR, offering a comprehensive overview that bridges foundational concepts with the forefront of innovation.
This survey builds upon the previous work Hitchhiker’s Guide to Super-Resolution [16], which gives a broad overview of the image SR field in general. We explore the foundational principles of DMs, dive into the challenges specific to SR, and examine the various types of DMs (DDPMs [4], SGMs [7], SDEs [17]), their relationship with other generative models, efficient sampling techniques, and applications in fields demanding high precision, such as medical imaging and remote sensing. Further, we investigate alternative input domains, architectural designs, and conditioning strategies and critically analyze current strengths and weaknesses.
@@
Concluding with a discussion on emerging trends and the potential they hold for reshaping SR and DM development, this survey sets the stage for future research. By offering clarity and direction in the rapidly evolving domain of DMs, we aim to inspire and inform the next wave of research, fostering advancements that continue to push the boundaries of what is possible in image SR with DMs.
@@
section VI - Domain-Specific Applications: DM-based SR applications, namely medical imaging, blind face restoration, atmospheric turbulence in face SR, and remote sensing.
section VII - Discussion and Future Work: Common problems of DMs for image SR and noteworthy research avenues for DMs specific to image SR.
section VIII - Conclusion: Summarizes the survey.
@@ II Image Super-Resolution
Image Super-Resolution (SR) represents a collection of methodologies designed to construct High-Resolution (HR) images from one or more Low-Resolution (LR) images. The domain of SR can be broadly categorized into two areas [16]: Single Image Super-Resolution (SISR) and Multi-Image Super-Resolution (MISR). In the SISR paradigm, a single LR image is processed to yield a single HR image. In contrast, MISR methods utilize multiple LR images to synthesize one or many HR outputs. This section primarily focuses on the SISR facet. It explores relevant datasets, established SR models, and techniques employed to assess image quality.
Given a LR image $\mathbf{x}\in\mathbb{R}^{\bar{w}\times\bar{h}\times c}$, the objective is to produce a corresponding HR image $\mathbf{y}\in\mathbb{R}^{w\times h\times c}$, where $\bar{w}<w$ and $\bar{h}<h$. The relationship is represented through a degradation mapping, defined as
$\mathbf{x}=\mathcal{D}\left(\mathbf{y};\Theta\right)$ (1)
where $\mathcal{D}$ is the degradation map $\mathcal{D}:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{\bar{w}\times\bar{h}\times c}$ and $\Theta$ signifies the degradation parameters, which can include aspects such as blur type, noise, scaling factor, among others.
The degradation is typically uncertain, posing the challenge for SISR to discern the inverse mapping of $\mathcal{D}$. The inherent complexity of this task arises from the ill-posed nature of predicting $\Theta$, leading to an optimization task aimed at minimizing the discrepancy between the predicted SR image $\mathbf{\hat{y}}$ and the original HR image $\mathbf{y}$:
$\hat{\Theta}=\operatorname{argmin}_{\Theta}\mathcal{L}\left(\mathbf{\hat{y}},\mathbf{y}\right)+\lambda\phi(\Theta),$ (2)
where $\mathcal{L}$ represents the loss between the predicted SR image and the actual HR image. Here, $\lambda$ is a balancing parameter, while $\phi(x)$ is introduced as a regularization term.
@@ II-A Datasets
Several datasets, offering a variety of images, resolutions, and content types, are available for SISR and instrumental in effectively training and evaluating SR models. Typically, these datasets consist of LR and HR image pairs. However, certain datasets contain only HR images, with the pairs created by downsampling them using techniques such as bicubic downsampling with anti-aliasing - a default setting for imresize in MATLAB [18]. For natural images, commonly employed datasets include the Berkeley Segmentation Dataset (BSDS300) [19], Flickr2K [20], and the Diverse 2K resolution (DIV2K) dataset [21]. Additionally, datasets such as ImageNet [22] and the Visual Object Classes (VOC2012) [23], which are favored in the Computer Vision (CV) community, are also utilized for SISR tasks. In face SR, datasets featuring high-quality, diverse images of human faces are required. To meet this demand, datasets like CelebA-HQ [24] and Flickr-Faces-HQ (FFHQ) [25] are commonly employed.
@@ II-B SR Models
where $\mathbf{\hat{y}}$ is the predicted HR approximation of the LR image $\mathbf{x}$ and $\theta$ the parameters of $\mathcal{M}$. Consequently, this translates into an optimization objective that minimizes Equation 2, loss function $\mathcal{L}$ between the estimation $\mathbf{\hat{y}}$ and the ground-truth HR image $\mathbf{y}$. The following sections will focus on standard methods to design a SR model, especially conventional deep learning methods, before we examine in detail how diffusion models fulfill the role of a SR model.
II-B1 Traditional Methods
@@
Traditional methods for image SR define a range of methodologies, such as statistical [26], edge-based [27], [28] patch-based [29], [30], prediction-based [31], [32] and sparse representation techniques [33]. They fundamentally rely on image statistics and the information inherent in existing pixels to generate HR images. Despite their utility, a noteworthy drawback of these methods is the potential introduction of noise, blur, and visual artifacts [16].
II-B2 Deep Learning Methods
Image SR has significantly evolved with advancements in Deep Learning (DL) and computational power. In general, DL-based approaches provide notable improvements over traditional methods. Typically, they employ a learning Convolutional Neural Network (CNN) for end-to-end mapping from LR to HR. Initial models, such as SRCNN [34], FSRCNN [35], and ESPCNN [36], utilized simple CNNs of diverse depth and feature maps sizes.
@@ II-B2 Deep Learning Methods
Later models adapted concepts from the broader CV domain into SR models, e.g., ResNet led to SRResNet, where residual information was propagated to successive network layers [37]. Likewise, DenseNet was adapted to image SR[38], e.g., SRDenseNet [39] and ESRGAN [40]. They employ dense blocks, where each layer receives additionally the features generated in all preceding layers. Recursive CNNs that recursively use the same module to learn representations were also inspired by other CV methods for DL-based SR methods in DRCN [41], DRRN [42] and CARN [43]. More recently, attention mechanisms have been incorporated to focus on regions of interest in images, predominantly via the channel and spatial attention mechanisms [16].
DL-based models can be further categorized into three directions based on the objective used to train the neural network: Pixel-based, GAN-based, and Flow-based techniques, which are introduced next.
II-B3 Pixel-based Methods
Pixel-based techniques utilize loss functions designed to minimize the difference directly in pixel space. Commonly used loss functions are the L1 and L2 losses. Models trained with pixel-based loss functions tend to generate images with a tendency toward blurriness and excessive smoothing [16]. This can result in less high-frequency details and even lead to unrealistic visual appearances. One way to overcome this limitation is to use generative models, such as GANs, flow-based methods, or diffusion models.
II-B4 Generative Adversarial Networks (GANs)
@@
GANs use two CNNs: A generator G and a discriminator D, which are trained simultaneously. The generator aims to produce HR samples that are as close to the original as to fool the discriminator, which tries to distinguish between generated and real samples. This framework, e.g., in SRGAN [44] or ESRGAN [40], is optimized using a combination of adversarial loss and content loss to produce less-smoothed images. The resultant images of state-of-the-art GANs are sharper and more detailed. Due to their capability to generate high-quality and diverse images, they have received much attention lately. However, they are susceptible to mode collapse, have a sizeable computational footprint, sometimes fail to converge, and require additional methods for stabilization [11].
II-B5 Flow-based Methods
@@
Flow-based methods employ optical flow algorithms to generate SR images [45]. They were introduced in an attempt to counter the ill-posed nature of image SR by learning the conditional distribution of plausible HR images given a LR input. They introduce a conditional normalized flow architecture that aligns LR and HR images by calculating the displacement field between them and then uses this information to recover SR images. They employ a fully invertible encoder capable of mapping any input HR image to the latent flow space and ensuring exact reconstruction. This framework enables the SR model to learn rich distributions using exact log-likelihood-based training [45]. This facilitates flow-based methods to circumvent training instability but incurs a substantial computational cost.
@@ II-C Image Quality Assessment (IQA)
Image quality is a multifaceted concept that addresses various properties, such as sharpness, contrast, and absence of noise. Hence, a fair evaluation of SR models based on produced image quality forms a non-trivial task. This section presents the essential methods, especially for diffusion models, to assess image quality in the context of image SR, which fall under the umbrella term Image Quality Assessment (IQA)  [^1] More SR-related IQA methods can be found in Moser et al. [16].. At its core, IQA refers to any metric that resembles the perceptual evaluations from human observers, specifically, the level of realism perceived in an image after the application of SR techniques. During this section, we will use the following notation: $N_{\mathbf{x}}=w\cdot h\cdot c$, which defines the number of pixels of an image $\mathbf{x}\in\mathbb{R}^{w\times h\times c}$ and $\Omega_{\mathbf{x}}=\{\left(i,j,k\right)\in\mathbb{N}_{1}^{3}|i\leq h,j\leq w,k\leq c\}$ that defines the set of all valid positions in $\mathbf{x}$.
II-C1 Peak Signal-to-Noise Ratio (PSNR)
The Peak Signal-to-Noise Ratio (PSNR) is one of the most widely used techniques to evaluate SISR reconstruction quality. It represents the ratio between the maximum pixel value L and the Mean Squared Error (MSE) between the SR image $\mathbf{\hat{y}}$ and the actual image $\mathbf{y}$.
@@ II-C1 Peak Signal-to-Noise Ratio (PSNR)
II-C2 Structural Similarity Index (SSIM)
The SSIM, like the PSNR, is a popular evaluation method that focuses on the differences in structural features between images. It independently captures the structural similarity by comparing luminance, contrast, and structures. SSIM estimates for an image $\mathbf{y}$ the luminance $\mu_{\mathbf{y}}$ as the mean of the intensity, while it is estimating contrast $\sigma_{\mathbf{y}}$ as its standard deviation:
@@ II-C2 Structural Similarity Index (SSIM)
II-C3 Mean Opinion Score (MOS)
The MOS is a subjective measure that leverages human perceptual quality for the evaluation of the generated SR images. Human viewers are shown SR images and asked to rate them with quality scores that are then mapped to numerical values and later averaged. Typically, these range from 1 (bad) to 5 (good) but may vary [15]. While this method is a direct evaluation of human perception, it is more time-consuming and cumbersome to conduct compared to objective metrics. Moreover, due to the highly subjective nature of this metric, it is susceptible to bias.
II-C4 Consistency
Consistency measures the degree of stability of non-deterministic SR methods, such as generative models like GANs or DMs. Like flow-based methods, generative approaches are intentionally designed to generate a spectrum of plausible outputs for the same input. However, low consistency is not desirable. Minor variations lessen the influence of a relatively consistent method in the input. Nevertheless, consistency can vary depending on the requirements. One commonly employed metric to quantify consistency is the Mean Squared Error.
II-C5 Learned Perceptual Image Patch Similarity (LPIPS)
Contrary to the pixel-based evaluation of PSNR and SSIM, the Learned Perceptual Image Patch Similarity (LPIPS) utilizes a pre-trained CNN $\varphi$, e.g., VGG [46] or AlexNet [47], and generates $L$ feature maps from the SR and HR image, and subsequently calculates the similarity between them. Given $h_{l}$ and $w_{l}$ as the height and width of the $l$-th feature map respectively, and a scaling vector $\alpha_{l}\in\mathbb{R}^{C_{l}}$, the LPIPS metric is formulated as follows:

### v3

@@ Abstract
Super-Resolution, Diffusion Models, Survey.
@@ I Introduction
In the ever-evolving field of computer vision, the task of image Super-Resolution (SR) – enhancing Low-Resolution (LR) images into High-Resolution (HR) counterparts – stands as a longstanding challenge due to its ill-posed nature: Multiple HR images are plausible for any given LR image, differing in aspects such as brightness and color [1]. Its applications span a broad spectrum, from everyday photography to refining satellite [2] and medical images [3]. Despite notable achievements of prior generative SR models, each comes with its own limitations. For example, the computational demands of Autoregressive models often outweigh their utility, while Normalizing Flows or Variational Autoencoders struggle to match quality expectations [4, 5, 6]. Although powerful, Generative Adversarial Networks (GANs) need careful regularization and optimization strategies to overcome instability issues [7].
@@
The advent of Diffusion Models (DMs) marks a significant shift in image generation tasks, including SR, challenging the long-standing dominance of Generative Adversarial Networks (GANs) [8, 9, 10, 11]. Applications like Dall-E and Stable Diffusion demonstrate that DMs have surpassed GANs in various aspects [12, 13, 14]. Their capability to generate high-quality images from LR inputs has shown immense promise in SR by closely aligning with the qualitative judgments of human evaluators [15]. In other words, human raters perceive SR images generated by DMs as more realistic than those produced by other generative models like GANs.
@@
However, as the volume of publications expands, staying updated on the latest developments is becoming more challenging, particularly for those new to the field. DMs diverge fundamentally from prior generative models and pose new challenges while addressing the limitations of earlier models. Identifying coherent trends and potential research directions is challenging despite this rapid expansion. This survey aims to demystify DMs, offers a comprehensive overview that bridges foundational concepts with the forefront of image SR and critically analyzes current strengths and weaknesses.
The presented survey builds upon the previous work Hitchhiker’s Guide to Super-Resolution [16], which gives a broad overview of the image SR field in general. Similar in spirit is the survey of Li et al., which reviews diffusion models on the more general image restoration tasks like inpainting and dehazing [17]. Both have overlapping topics, such as the foundations and types of DMs, namely DDPMs [8], SGMs [11], and SDEs [10]. Moreover, both surveys highlight the introduction of conditioning strategies and zero-shot diffusion as well as show potential research directions. However, our survey covers all topics related to image SR and is, therefore, more detailed regarding recent developments specifically developed for image SR. Moreover, we explain SR-related challenges, like color shifting and cascaded image SR. We also highlight the relationship of DMs with other generative SR models, namely Variational Autoencoders, GANs, and Flow-based methods. In addition, we review frequency-based DMs, alternative corruption spaces, and diffusion-based image SR applications.
@@
Concluding with a discussion on emerging trends and their potential for reshaping SR and DM development, this survey sets the stage for future research. By offering clarity and direction in the rapidly evolving domain of DMs, we aim to inspire and inform the next wave of research, fostering advancements that continue to push the boundaries of what is possible in image SR with DMs.
@@
section VII - Domain-Specific Applications: DM-based SR applications, namely medical imaging, blind face restoration, atmospheric turbulence in face SR, and remote sensing.
section VIII - Discussion and Future Work: Common problems of DMs for image SR and noteworthy research avenues for DMs specific to image SR.
section IX - Conclusion: Summarizes the survey.
@@ II Image Super-Resolution
The goal of image Super-Resolution (SR) is to transform one or more Low-Resolution (LR) images into High-Resolution (HR) images. The domain can be broadly categorized into two areas [16]: Single Image Super-Resolution (SISR) and Multi-Image Super-Resolution (MISR). In SISR, a single LR image leads to a single HR image. In contrast, MISR methods use multiple LR images to produce one or many HR outputs. This section focuses on SISR and explores relevant datasets, established SR models, and techniques to assess image quality.
Given a LR image $\mathbf{x}\in\mathbb{R}^{\bar{w}\times\bar{h}\times c}$, the goal is to generate a HR counterpart $\mathbf{y}\in\mathbb{R}^{w\times h\times c}$ with $\bar{w}<w$ and $\bar{h}<h$. The relationship is represented by a degradation mapping
$\mathbf{x}=\mathcal{D}\left(\mathbf{y};\Theta\right)=((\mathbf{y}\otimes\boldsymbol{k})\downarrow_{s}+\boldsymbol{n})_{JPEG_{q}}$ (1)
where $\mathcal{D}$ is a degradation map $\mathcal{D}:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{\bar{w}\times\bar{h}\times c}$ and $\Theta$ contains degradation parameters, including aspects like blur $\boldsymbol{k}$, noise $\boldsymbol{n}$, scaling $s$, and compression quality $q$ [18]. The degradation is typically unknown, posing the main challenge in determining the inverse mapping of $\mathcal{D}$ with parameters $\theta$, usually embodied as SR model. It leads to an optimization task aimed at minimizing the difference between the predicted SR image $\mathbf{\hat{y}}$ and the original HR image $\mathbf{y}$:
$\theta^{*}=\operatorname{argmin}_{\theta}\mathcal{L}\left(\mathbf{\hat{y}},\mathbf{y}\right)+\lambda\phi(\theta),$ (2)
where $\mathcal{L}$ represents the loss between the predicted SR image and the actual HR image. Here, $\lambda$ is a balancing parameter, while $\phi(\theta)$ is introduced as a regularization term.
The inherent complexity arises from the ill-posed nature of predicting $\theta$, as several SR images can be valid for any given LR image, i.e., they can have similar loss values compared to the ground-truth image but are subjectively perceived differently due to many aspects like brightness and coloring [19, 1, 15]. Traditional regression techniques, like standard CNNs, are often adequate for lower magnifications but struggle to replicate high-frequency details required at higher magnifications (e.g., $s>4$). To address this, SR models must hallucinate realistic details beyond interpolation, which typically falls under the umbrella topic of generative models, where diffusion models are now at the forefront.
@@ II-A Datasets
Several datasets offer a variety of images, resolutions, and content types. Typically, these datasets consist of LR and HR image pairs. However, some datasets contain only HR images, with LR images created by bicubic downsampling with anti-aliasing - a default setting for imresize in MATLAB [20]. One famous general SR train set is the Diverse 2K resolution (DIV2K) dataset [21], which includes various realistic images at different resolutions designed specifically for image SR. Classical test datasets for SR models trained on DIV2K are Set5 [22], Set14 [23], BSDS100 [24], Urban100 [25] and Manga109 [26] that cover a variety of scenes and images contents like buildings and manga paintings. Flickr2K [27] and Flickr-Faces-HQ (FFHQ) [28] offer diverse sets of human-centric and scene-centric images from Flickr, respectively. While FFHQ is commonly employed for training models for face SR tasks, Flickr2K is usually used as a train data extension in combination with DIV2K. Another dataset for face SR is CelebA-HQ [29], which provides high-quality celebrity images and is typically used to evaluate FFHQ-trained SR models. For broader applications in CV, datasets like ImageNet [30] and Visual Object Classes (VOC2012) [31] are favored. ImageNet offers an extensive range of images that help train models on various object classes, whereas VOC2012 is vital for object detection and segmentation. Both are valuable for multi-task learning involving SR. More datasets can be found in the Hitchhiker’s Guide to Super-Resolution [16].
@@ II-B SR Models
where $\mathbf{\hat{y}}$ is the predicted HR approximation of the LR image $\mathbf{x}$ and $\theta$ the parameters of $\mathcal{M}$. The parameters $\theta$ are optimized using Equation 2, i.e., minimizing the loss function $\mathcal{L}$ between the estimation $\mathbf{\hat{y}}$ and the ground-truth HR image $\mathbf{y}$. The following section focuses on standard methods for designing an SR model, especially deep learning methods before we examine how diffusion models fulfill this role in detail.
@@
Traditional Methods: Traditional methods for image SR define a range of methodologies, such as statistical [32], edge-based [33], [34] patch-based [35], [36], prediction-based [37], [38] and sparse representation techniques [39]. They fundamentally rely on image statistics and the information inherent in existing pixels to generate HR images. Despite their utility, a noteworthy drawback of these methods is the potential introduction of noise, blur, and visual artifacts [16].
@@
Regression-based Deep Learning: Image SR significantly evolved with advancements in deep learning and computational power. Typically, they employ a Convolutional Neural Network (CNN) for end-to-end mapping from LR to HR. Initial models, such as SRCNN [40], FSRCNN [41], and ESPCNN [42], utilized simple CNNs of diverse depth and feature maps sizes. Later models adapted concepts from the broader CV domain into SR models, e.g., ResNet led to SRResNet, where residual information was propagated to successive network layers [43]. Likewise, DenseNet[44] was adapted with SRDenseNet [45]. They employ dense blocks, where each layer receives additionally the features generated in all preceding layers. Recursive CNNs that recursively use the same module to learn representations were also inspired by other CV methods for regression-based SR methods in DRCN [46], DRRN [47], and CARN [48]. More recently, attention mechanisms have been incorporated to focus on regions of interest in images, predominantly via the channel and spatial attention mechanisms [16, 49, 50, 51]. All those methods have in common that they are regression-based. Commonly used loss functions are the L1 and L2 losses. As mentioned, they often produce satisfying results for lower magnifications but struggle to replicate the high-frequency details required at higher magnifications (e.g., $s>4$). These limitations arise because these models primarily learn an averaged mapping (due to L1 and L2 losses) from LR to HR images, which tends to produce overly smooth textures lacking detail, especially noticeable in larger upscaling factors [16]. To address this, SR models must hallucinate realistic details beyond simple interpolation, a challenge typically tackled by generative models.
@@
Generative Adversarial Networks (GANs): One of the most prominent generative models is the Generative Adversarial Network (GAN). It uses two CNNs: A generator G and a discriminator D, which are trained simultaneously. The generator aims to produce HR samples that are as close to the original as to fool the discriminator, which tries to distinguish between generated and real samples. This framework, e.g., in SRGAN [52] or ESRGAN [53], is optimized using a combination of adversarial loss and content loss to produce less-smoothed images. The resultant images of state-of-the-art GANs are sharper and more detailed. Due to their capability to generate high-quality and diverse images, they have received much attention lately. However, they are susceptible to mode collapse, have a sizeable computational footprint, sometimes fail to converge, and suffer from stabilization issues [7].
@@
Flow-based Methods: Flow-based methods employ optical flow algorithms to generate SR images [54]. They were introduced in an attempt to counter the ill-posed nature of image SR by learning the conditional distribution of plausible HR images given a LR input. They introduce a conditional normalized flow architecture that aligns LR and HR images by calculating the displacement field between them and then uses this information to recover SR images. They employ a fully invertible encoder capable of mapping any input HR image to the latent flow space and ensuring exact reconstruction. This framework enables the SR model to learn rich distributions using exact log-likelihood-based training [54]. This facilitates flow-based methods to circumvent training instability but incurs a substantial computational cost.
@@ II-C Image Quality Assessment (IQA)
Image quality is a multifaceted concept that addresses properties like sharpness, contrast, and absence of noise. Hence, a fair evaluation of SR models based on produced image quality forms a non-trivial task. This section presents the essential methods, especially for diffusion models, to assess image quality in the context of image SR, which fall under the umbrella term Image Quality Assessment (IQA)  [^1] More SR-related IQA methods can be found in Moser et al. [16].. At its core, IQA refers to any metric that resembles the perceptual evaluations from human observers, specifically, the level of realism perceived in an image after the application of SR techniques. During this section, we will use the following notation: $N_{\mathbf{x}}=w\cdot h\cdot c$, which defines the number of pixels of an image $\mathbf{x}\in\mathbb{R}^{w\times h\times c}$ and $\Omega_{\mathbf{x}}=\{\left(i,j,k\right)\in\mathbb{N}_{1}^{3}|i\leq h,j\leq w,k\leq c\}$ that defines the set of all valid positions in $\mathbf{x}$.
Peak Signal-to-Noise Ratio (PSNR): The Peak Signal-to-Noise Ratio (PSNR) is one of the most widely used techniques to evaluate SISR reconstruction quality. It represents the ratio between the maximum pixel value L and the Mean Squared Error (MSE) between the SR image $\mathbf{\hat{y}}$ and the HR image $\mathbf{y}$.
@@
Structural Similarity Index (SSIM): The SSIM, like the PSNR, is a popular evaluation method that focuses on the differences in structural features between images. It independently captures the structural similarity by comparing luminance, contrast, and structures. SSIM estimates for an image $\mathbf{y}$ the luminance $\mu_{\mathbf{y}}$ as the mean of the intensity, while it is estimating contrast $\sigma_{\mathbf{y}}$ as its standard deviation:
@@
Mean Opinion Score (MOS): The MOS is a subjective measure that leverages human perceptual quality for the evaluation of the generated SR images. Human viewers are shown SR images and asked to rate them with quality scores that are then mapped to numerical values and later averaged. Typically, these range from 1 (bad) to 5 (good) but may vary [15]. While this method is a direct evaluation of human perception, it is more time-consuming and cumbersome to conduct compared to objective metrics. Moreover, due to the highly subjective nature of this metric, it is susceptible to bias.
Consistency: Consistency measures the degree of stability of non-deterministic SR methods, such as generative models like GANs or DMs. Like flow-based methods, generative approaches are intentionally designed to generate a spectrum of plausible outputs for the same input. However, low consistency is not desirable. Minor variations lessen the influence of a relatively consistent method in the input. Nevertheless, consistency can vary depending on the requirements. One commonly employed metric to quantify consistency is the Mean Squared Error.
Learned Perceptual Image Patch Similarity (LPIPS): Contrary to the pixel-based evaluation of PSNR and SSIM, the Learned Perceptual Image Patch Similarity (LPIPS) utilizes a pre-trained CNN $\varphi$, e.g., VGG [55] or AlexNet [56], and generates $L$ feature maps from the SR and HR image, and subsequently calculates the similarity between them. Given $h_{l}$ and $w_{l}$ as the height and width of the $l$-th feature map respectively, and a scaling vector $\alpha_{l}\in\mathbb{R}^{C_{l}}$, the LPIPS metric is formulated as follows:
@@
No-Reference Metrics: All IQA metrics discussed so far require a reference (ground-truth) image. However, there are cases where no reference images are available, e.g., in unsupervised settings. Fortunately, we can assess an image by measuring the distance of statistical features from those obtained from a collection of high-quality images of a similar domain, i.e., natural images. This can be opinion- and distortion-aware like BRISQUE [57] or opinion- and distortion-unaware like NIQE [58]. Another intriguing way to assess no-reference image quality is to exploit the visual-language pre-trained CLIP model [59]. One example is CLIP-IQA, which calculates the cosine similarity of the encoded image with two prompts of opposing meaning, i.e., ”good photo” and ”bad photo” [60]. The resulting relative similarity metric for one or the other prompt determines the image quality. CLIP-IQA shows results comparable to those of BRISQUE without the hand-crafted features and surpasses other no-reference IQA methods like NIQE. Another way to exploit deep learning models is to train them to predict subjective scores using IQA datasets like TID2013 [61]. Examples are DeepQA [62], NIMA [63], or MUSIQ [64]. Others can be found in the learning-based perceptual quality section of the Hitchhiker’s Guide to Super-Resolution [16].